OpenStudy (anonymous):

1/3 (2x - 5) = 3/4 (x-2)

OpenStudy (turingtest):

is it (1/3)(2x-5)=(3/4)(x-2) ?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

multiply both sides by their LCD which is 3x4=12: 4(2x-5)=9(x-2) 8x-20=9x-18 combine like terms: -20+18=9x-8x -2=x